OpenStudy (luigi0210):

Related rates problems:

OpenStudy (luigi0210):

A pebble is dropped into a calm pond, causing ripples in the form of concentric circles. The radius \(\large r\) of the outer circle is increasing at \(\large 1 ft/sec\) When the radius equals \(\LARGE 4ft\), at what rate is the total area of the disturbed water changing?

OpenStudy (anonymous):

rate of change of radius =dr/dt=1ft/sec

OpenStudy (anonymous):

now area of circle=A=pi*r^2 rate of change of area=dA/dt=pi*2*r*dr/dt at r=4feet dA/dt=pi*2*4*1.........as dr/dt=1 dA/dt=8*pi

OpenStudy (anonymous):

=rate of change of area at r=4

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Hmm, makes sense, thanks :)

OpenStudy (anonymous):

u got it right ??

OpenStudy (anonymous):

im to stupid for this D: thanks for helping me to

OpenStudy (luigi0210):

Yea I got it :)